# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v6

@@ B.6 Estimating Average Head’s Monthly Income
[…]
@@
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$. Source: Created by the author using the THBS dataset.
Figure B.6: Distribution of the THBS Heads’ Age
The first issue is that the wages of factory workers are aggregated in a “15 years old or older” category in the ASCT (volume 17). Figure B.6 shows the THBS heads’ age distribution. The sample mean is $33.2$ years old, which is unchanged if I exclude a head aged more than $60$ (i.e., $32.1$ years old). This means that I have to refer to the average daily wage for the 30s to 40s for comparison. However, since many workers are in their 20s in the distribution of entire workers, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old. Thus, I need the wage profiles by age to correct for this downward bias.888888 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages. As is clear, the same problem arises in this case.
@@ B.6.1 Daily Wage Distribution by Ages
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30-39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$. Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
@@
The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$. Since $m=2$, there are ${}_{8}\mathrm{H}_{2}$ ($=36$) repeated combinations for the candidates of the parameters. By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood. The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$. [^89] The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$. Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models. This statistic has an asymptotic chi-squared distribution with $m$ degree of freedom. The estimated deviance difference with the simplest linear model (i.e., $m=1$ and $p_{1}=1$) is $54.6$ with $p$-value = $0.000$. Similarly, the difference with an another non-linear model with $m=1$ and $p_{1}=-2$ is $34.2$ with $p$-value = $0.000$. The projection model fitted using the best power vector is as follows:
@@
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference. [^90] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.919191This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen. In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged 15 and older is 1.14, while the overall average is 1.13, a difference of only 0.01 yen. At least, in the light of the first moment, this confirms that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.929292 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@
(a) 1918
(b) 1919
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12-14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25-65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP. Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
Figure B.8: Daily Wage by Ages Distribution:
Shifted Fractional Polynomial
@@ B.6.2 Ancillary Wage
Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage. The raw data are listed in Table B.3. The dashed line indicates the linear fit using the log-transformed variables: the fitted regression is $\hat{y}=0.747+0.428x$ with $R^{2}=0.82$. The predictions are materially similar if I use the fractional polynomial using the raw data on the number of factory workers and share of ancillary wage in Table B.3. In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$. I prefer to use linear regression herein because it can provide more precise fitted values for factories with a small number of workers. Source: Kitazawa (1924, pp. 20; 58–62).
@@
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is 12%, which is slightly larger than the value for the entire city of Tokyo (9%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. [^93] Note again that the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, say the Tsukishima factories but not to calculate the average ancillary wage for the THBS sample itself.
@@
Notes: Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21). Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively. Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412). Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). Data on the number of heads are from the THBS sample.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
Table B.4: Estimating the Monthly Income of Adult Male Factory Workers in Machinery Sector
@@
$n_{8}$ 10 10
$n_{9}$ 6 6
$n_{10}$ 5 5
[…]
@@
[…]
Panel A in Table B.4 summarizes all the components in equation 14. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $55$ yen, which is approximately $2$ yen higher than the average monthly earnings in the THBS heads ($53$ yen, see Panel C of Table 2) as I expected in Online Appendix B.6.2. Therefore, the deviation from the population mean ($55$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $2$ yen difference is less than one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $53.6$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. When I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).

### v7

@@ B.6 Estimating Average Head’s Monthly Income
$\textit{Wage}^{15+}_{Machine}=\frac{\sum_{c}\{(\textit{Daily Wage}^{15+}_{Machine,c}\times\textit{Working Days}_{Machine,c})\times\textit{Workers}^{15+}_{Machine,c}\}}{12\sum_{c}\textit{Workers}^{15+}_{Machine,c}}.$ (5)
@@
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$.
Source: Created by the author using the THBS dataset.
The first issue is that the wages of factory workers are aggregated in a “15 years old or older” category in the ASCT (volume 17). Figure B.6 shows the THBS heads’ age distribution. The sample mean is $33.2$ years old, which is unchanged if I exclude a head aged more than $60$ (i.e., $32.1$ years old). This means that I have to refer to the average daily wage for the 30s to 40s for comparison. However, since many workers are in their 20s in the distribution of entire workers, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old. Thus, I need the wage profiles by age to correct for this downward bias.101101101 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages. As is clear, the same problem arises in this case.
@@ B.6.1 Daily Wage Distribution by Ages
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30-39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$.
Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
@@
The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$. Since $m=2$, there are ${}_{8}\mathrm{H}_{2}$ ($=36$) repeated combinations for the candidates of the parameters. By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood. The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$.102102102The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$. Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models. This statistic has an asymptotic chi-squared distribution with $m$ degree of freedom. The estimated deviance difference with the simplest linear model (i.e., $m=1$ and $p_{1}=1$) is $54.6$ with $p$-value = $0.000$. Similarly, the difference with an another non-linear model with $m=1$ and $p_{1}=-2$ is $34.2$ with $p$-value = $0.000$. The projection model fitted using the best power vector is as follows:
@@
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference.103103103For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.104104104This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen. In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged 15 and older is 1.14, while the overall average is 1.13, a difference of only 0.01 yen. At least, in the light of the first moment, this confirms that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.105105105 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12-14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25-65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP.
Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
@@ B.6.2 Ancillary Wage
Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage. The raw data are listed in Table B.3. The dashed line indicates the linear fit using the log-transformed variables: the fitted regression is $\hat{y}=0.747+0.428x$ with $R^{2}=0.82$. The predictions are materially similar if I use the fractional polynomial using the raw data on the number of factory workers and share of ancillary wage in Table B.3. In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$. I prefer to use linear regression herein because it can provide more precise fitted values for factories with a small number of workers.
Source: Kitazawa (1924, pp. 20; 58–62).
@@
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is 12%, which is slightly larger than the value for the entire city of Tokyo (9%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. Note again that, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, say the Tsukishima factories but not to calculate the average ancillary wage for the THBS sample itself.
@@
Notes:
Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21).
Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively.
Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412).
Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). Data on the number of heads are from the THBS sample.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
Table B.4: Estimating the Monthly Income of Adult Male
Factory Workers in Machinery Sector
@@
$n_{8}$ 8 8
$n_{9}$ 5 5
$n_{10}$ 4 4
Panel B: Estimated monthly earning: $w^{\textit{Total}}$
$w^{\textit{Total}}$ 56.0 54.5
Notes:
1. Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30-40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST. Details in the calculation steps are summarized in Online Appendix B.6.1.
2. Average annual working days ($l_{j}$): The ASCT documents the number of factories and average annual working days for each machinery sub-sector in 2 factory-size bins. I used the number of factories in each machinery sub-sector and factory size bin as the weights to calculate the weighted average of the annual working days in the entire machinery sector.
3. Weight for the ancillary wage ($\vartheta$): The average share of monthly ancillary wage in the machinery sector is calculated based on the factory survey report in Kitazawa (1924). The calculation steps are described in Online Appendix B.6.2. Column 1 uses the share based on the factory size distribution in Tsukishima, whereas column 2 uses that in the entire Tokyo City.
4. Frequency weights ($n_{m}$) are based on the number of cross-sectional observations between January and October 1919 in the adjusted monthly panel dataset.
5. $w^{\textit{Total}}$ is calculated using equation 14.
Sources: Data used to calculate the average daily wage are from the Tokyo City Office (1921, pp. 726-757); Tokyo City Office (1922, pp. 732-733); Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data used to calculate the average annual working days are from the Tokyo City Office (1921, pp. 726-741) and Tokyo City Office (1922, pp. 716-725). Data used to calculate the average share of monthly ancillary wage are from Kitazawa (1924, pp. 20; 58-62) and the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674).
Following the institution described in Online Appendix B.6.2, the monthly wage of the male workers in the skilled age range ($w^{\textit{Total}}$) can be defined as follows:
$w^{\textit{Total}}=w^{\textit{Regular}}+w^{\textit{Ancillary}},$ (9)
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.106106106Strictly, this weight is not required to be in this range. However, all the cases reported in Kitazawa (1924) take the weights within this range. The monthly wage equation is then specified as follows:
$w^{\textit{Total}}=(1+\vartheta)w^{\textit{Regular}}.$ (10)
To calculate equation 10, I have to conduct a few final tunings in my wage data. First, I fixed the deviations in the timing of the manufacturing censuses and the THBS observations. Let $w^{\textit{Regular}}_{j}$ be the monthly regular wage calculated using wage statistics measured in year $j\in\{1918,1919\}$ as follows:
$w^{\textit{Regular}}_{j}=w^{\textit{Daily}}_{j}l_{j}/12,$ (11)
where $w^{\textit{Daily}}_{j}$ is the average daily wage of male factory workers in the skilled age range and $l_{j}$ is the average annual working days (Online Appendix B.6.1). The THBS were conducted from November 1918 to January 1920, and my THBS sample was observed from December 1918 to December 1919 (Section 3). However, $w^{\textit{Regular}}_{j}$ is calculated using the wage statistics measured in the manufacturing census conducted in December of 1918 and 1919. This means that if I use $w^{\textit{Regular}}_{1919}$ alone for equation 10, it shall provide upward-biased wage due to the deviation between the timing of the THBS observations and of the manufacturing census. To cope with this issue, I use the weighted average between $w^{\textit{Regular}}_{1918}$ and $w^{\textit{Regular}}_{1919}$ to yield monthly wage series as follows:
$w^{\textit{Regular}}_{m}=\{(12-m)w^{\textit{Regular}}_{1918}+mw^{\textit{Regular}}_{1919}\}/12,$ (12)
@@
$w^{\textit{Total}}=(1+\vartheta)\frac{\sum_{m}n_{m}\{(12-m)w^{\textit{Daily}}_{1918}l_{1918}+mw^{\textit{Daily}}_{1919}l_{1919}\}}{12^{2}\sum_{m}n_{m}}.$ (14)
Panel A in Table B.4 summarizes all the components in equation 14. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $56$ yen, which is similar to the average and median of monthly earnings in the THBS heads ($59$ and $56$ yen, respectively; see Panel C of Table 2). Therefore, I can conclude that the deviation from the population mean ($56$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3$ yen difference from the average monthly income is roughly one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $54.5$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. If I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).